# The Impact of COVID-19 on Co-authorship and Economics Scholars’ Productivity

### v1

@@ Abstract
The COVID-19 pandemic has disrupted traditional academic collaboration patterns, prompting a unique opportunity to analyze the influence of peer effects and coauthorship dynamics on research output. Using a novel dataset, this paper endeavors to make a first cut at investigating the role of peer effects on the productivity of economics scholars, measured by the number of publications, in both pre-pandemic and pandemic times. Results show that peer effect is significant for the pre-pandemic time but not for the pandemic time. The findings contribute to our understanding of how research collaboration influences knowledge production and may help guide policies aimed at fostering collaboration and enhancing research productivity in the academic community.
Keywords COVID-19  $\cdot$ Productivity  $\cdot$ Co-authorship  $\cdot$ Game Theory  $\cdot$ Publication
@@ 1 Introduction
Academic productivity is commonly used in academic hiring and promotion decisions, as well as in evaluations of research programs and departments [1, 2], because a higher number of publications may indicate that the scholar has been active in research and has made significant contributions to the field. In recent years, there has been a growing interest in understanding the factors that contribute to the productivity of scholars in various disciplines. Among these factors, the role of peer effects and co-authorship has emerged as an area of significant importance [3]. Scholars who have strong social networks and interactions with colleagues in their field may publish more, because they have access to more resources, including research funding, data, and other research materials. Co-authorship also facilitates knowledge sharing, which could lead to new research ideas and opportunities for publication. For example, [4] demonstrates that one single strong connection could have a substantial and positive influence on the scholar’s productivity and citation rates.
@@
The COVID-19 pandemic has fundamentally altered the way we live, work, and collaborate. As the pandemic forced the closure of universities and research institutions, it inadvertently reshaped the landscape of academic research. Consequently, this unprecedented situation presents a unique opportunity to examine the peer effects and co-authorship dynamics among economics scholars in both pre-pandemic and pandemic times. Working at home may reshape scholars’ way of writting, communication, and collaboration. On one hand, scholars have more flexible schedules and greater autonomy over their work. It provides opportunities for virtual collaboration through video conferencing and instant messaging among scholars who locate in different regions or time zones. On the other hand, scholars who work from home may experience greater isolation and have less opportunity for informal interactions and discussions with their co-authors. It may also create distractions and disruptions, such as balancing work responsibilities with home responsibilities.
@@
This paper employs a novel dataset, collected from Google Scholar, that includes scholarly literature and co-authorship networks of economics scholars. We empirically estimate the peer effect of economic scholars on their number of publications during both the pre-pandemic and pandemic periods. The present study is important for several reasons. First, understanding the role of co-authorship in academic productivity provides insights into how research collaborations influence knowledge production and dissemination. Second, by comparing pre-pandemic and pandemic times, this study offers a better understanding of how the COVID-19 crisis has affected academic collaboration patterns and the overall productivity of scholars in the field of economics. Finally, the findings of this study may help guide policies aimed at fostering collaboration and enhancing research productivity in the academic community.
@@
This paper is organized as follows. Section 2 reviews the relevant literature on peer effects and academic productivity. Section 3 describes the data and provides summary statistics. Section 4 presents the model framework and empirical strategy, while Section 5 discusses the main results. Finally, Section 6 concludes and offers directions for future research.
@@ 2 Literature Reivew
From the methodological perspective, our work is closely related to two main strands of the literature. Firstly, fast-growing literature on peer effects in networks. As reflected by [5], there are two different impacts from peers. One is endogenous peer effects which are the impact of peers’ outcomes and the other are contextual peer effects, which are the impact of peers’ characteristics. Distinguishing these impacts may be impossible because of the simultaneity in the behavior of interacting agents. [6, 7, 8, 9] are the initial studies of network interactions. For example, [6] builds the benchmark linear-in-means model of peer effects and describes the identification conditions when agents interact through a network assuming peers of peers are not peers. The assumption that the agents’ characteristics have an impact on individual outcomes only through their effect on peers’ outcomes, provides valid instruments addressing correlated effects. Therefore, endogenous and contextual peer effects are identified. An important insight is that identification depends on the structure of the network itself. [10, 11, 12, 13] investigate heterogeneous peer effects, in which men and women are subject to different peer effects. Individuals could also be subject to different effects from male peers and from female peers, which potentially leads to endogenous peer effects. Concerning heterogeneity of peer effects, [14] incorporates heterogeneity analysis by assuming that endogenous peer effect coefficients are random in a linear-in-means model. He shows these random endogenous peer effects can be point-identified if there is no contextual peer effect for an exogenous characteristic.
@@
The second strand of literature is related to tackling the problem of correlated effects and exploiting the identification possibilities generated by interaction networks. In the literature, researchers have developed at least four broad strategies to address this correlated effects issue: random peers, random shocks, structural endogeneity, and panel data. The first strategy is random peers, who are randomly allocated through natural or designed experiments. For example, [15] looks at Dartmouth College roommates in pairs, triples, or quads among students. [16] randomly match workers in pairs in the lab. [7] look at the choice of a major among Bocconi undergraduates. The insight is with random peers an agent’s observed and unobserved characteristics are uncorrelated with their peers’ observed and unobserved characteristics. The second strategy is random shocks. For instance, [17] uses exogenous variations to study treatment randomization which allows researchers to identify the causal impacts of the treatment and peers’ treatments and peers’ outcomes even when the network is endogenous within a linear-in-means framework. [18, 19, 20] study spillover effects. As an individual’s potential outcome may depend on the full vector of potential treatments, the causal impact of a randomized treatment cannot be estimated by simply computing the difference in average outcome among treated and untreated individuals. [21, 22] identify spillovers by assuming agents are organized in groups and that spillovers take place within, not between groups, then comparing the outcomes of untreated individuals in treated and untreated groups. The third strategy is called the structural framework. [23] first proposed the structural approach to address correlated effects issues. This approach provides a potentially powerful way to control for network endogeneity in peer effect regressions, reminiscent of Heckman’s correction for sample network formation simultaneously may allow researchers to recover information on common unobservables. The fourth strategy on peer effects in networks applies to panel data. However, this literature is scarce. A few papers analyze peer effects utilizing panel data, see [24, 25, 26]. These studies have introduced individual fixed effects but do not address contextual peer effects associated with time-invariant characteristics. We view this as a potentially important limitation of these frameworks and thus is an important future research question.
@@
The model that is employed in the present study belongs to the structural endogeneity framework. It relates to the models that use static games with incomplete information in which, agents act non-cooperatively, see [27, 28]. The assumption of incomplete information of the peer effect models for discrete outcomes is broadly studied, for instance, [29, 30, 31, 32]. In the literature, agent $i$’s decision is influenced by their own observable characteristics, unobservable individual type, and other agents’ choice. The recent study from [33, 34] propose methods to estimate the network’s probability distribution using cross-sectional data when the network is imperfectly observed. They construct a network game, and each agent chooses an integer outcome to maximize his or her preference, which contains observed characteristics of the agent and the peers, the difference between the choice of the agent and the peers, a cost function, and a private signal. They prove that under a few assumptions, there is a unique Bayesian Nash Equilibrium for this game, and an estimator is proposed based on pseudo-likelihood maximization.
@@
The literature of empirical papers exploring the relationship between co-authorship and academic productivity has expanded recently. Nonetheless, consensus remains elusive regarding whether this relationship is positive, negative, or insignificant. For example, [35] demonstrates that economists who are more collaborative are also more productive. Factors such as tenure, age, and geographical variables do not have a significant impact on productivity. [36] also finds a positive correlation between intellectual collaboration and individual performance, after accounting for endogenous network formation, unobservable heterogeneity, and factors that vary over time. As indirect evidence, [37] identifies that, at the individual level, the average publication quality rises with the average number of authors per paper, individual field diversity, the total number of published papers, and the presence of foreign co-authors. Female and older academics tend to publish less frequently. Conversely, some scholars argue that the relationship between co-authorship and productivity exhibits a negative correlation or lacks statistical significance. [38] discovers that for a specific scholar, increased co-authorship leads to higher quality, longer, and more frequent publications. However, after adjusting for the number of authors, the relationship between co-authorship and a scholar’s attributable output becomes negative. Other research contends that after controlling for article length, journal and author quality, and subject area, scholar fixed effects, the productivity of prior collaborators is not a significant determinant of a researcher’s own productivity [39], or higher quality research [40]. [41] finds that coauthors of highly helpful scientists that die experience a decrease in output quality but not output quantity.
@@ 3 Economic Scholars Data
The Economic scholars dataset consists of 1,671 core faculties from the best 50 Economics Schools in the United States based on US News in 2022, who have registered themselves a homepage on Google Scholar. The dataset does not include visiting professors, teaching professors, or lecturers. The homepage provides rich information on the individual’s influences and research journey in academia. Besides the names and affiliations, scholars may list their research interests and sub-fields at the head of the page. On the right-hand side, the scholar’s cumulative citations, H-index, and I10 index are shown, along with regular coauthors and the histogram of the number of citations and publications each year for the last 10 years. On the left-hand side, a comprehensive list of academic output of the scholar’s could be viewed, including the paper’s title, journal, authors, year, and the number of citations. See a snapshot of the Google Scholar homepage in the figure below.
@@
A key feature of the data collected from Google Scholar is its comprehensiveness. The website aggregates information from a wide range of sources, including journal articles, conference papers, theses and dissertations, technical reports, books and book chapters, patents, working papers, and repositories. Thus compared to using data sources from academic journals, the lagged effect of co-authorship on scholars’ academic output due to journals’ reviewing process could be largely alleviated.
Based on scholars’ publications, we construct an economics scholar co-authorship network each year for the last 10 years. Each node represents a faculty member in the economics department, and there is an edge between any two nodes if they have co-authored at least one paper in a given year. The edge density of a temporal network indicates the prevalence of scholar collaborations in that year, or the percentage of observed co-authorships over all possible collaborations between any two scholars. Economic scholars’ productivity, revealed by the annual average numbers of publications, gradually increase since 2013, experience a surge and reaches the top in 2020, then falls back slowly in 2021 and 2022. Network edge density reaches its highest level in 2018, then declines and reaches the bottom in 2020, before it bounces back in 2021 and 2022. Interestingly, scholars’ average productivity moves in the same direction as the prevalence of collaborations before 2018, but in the opposite direction from 2019 to 2021.
@@
The experience of economic faculties in academia is 25 years on average as of 2023, which is measured by the difference between 2022 and the year of the scholar’s first published article on Google Scholar. African American scholars, identified either directly by their listed nationality in CV or indirectly by the combination of the predicted country of origin using names and faculty pictures, account for 0.9% of the economics faculty. 18% of the faculty in the sample are female. We also label the expertise of economic scholars with their listed sub-fields on the Google homepage. For those who have left the section blank, We fill in the missing values by the fields that are shown on the department faculty website and the scholar’s curriculum vitae. The sub-fields are not mutually exclusive, and it is reasonable for scholars to have two, three, or more specialties. For example, a macroeconomist may specify econometrics, monetary policy, international economics, or finance as the expertise as well. In our sample, economic/econometrics theory, macroeconomics, and labor economics are the three top fields that scholars may work on, followed by econometrics, industrial organization, and development economics.
Table 1: Summary Statistics of Economics Scholars
@@
Table 2: Sub-fields of Economics Scholars
@@ 4 Network Game with Peer Effects and Incomplete Information
Following the notation in [34], suppose there is a population of $n$ agents that interact through a network matrix $\mathbf{G}$ with zero diagonal and non-negative elements $g_{ij}$ that represents the proximity of $i$ and $j$. Each agent $i$ chooses an integer outcome, $y_{i}$, to maximize his or her individual utility:
@@
Define $\Delta$ as the first difference operator. It is proved in [34] that under the following three assumptions, there is a unique integer $r_{0}$ that maximizes the preference $U_{i}^{e}\left(\cdot\right)$, and $U_{i}^{e}\left(r\right)\geq\max\left\{U_{i}^{e}\left(r-1\right),U_{i}^{e}\left(r+1\right)\right\}$ if and only if $r=r_{0}$:
@@
where $\Phi(\cdot)$ is the CDF of standard normal distribution, $\mathbf{L}$ is a mapping, $\mathbf{y}^{e}=\mathbf{L}(\mathbf{\theta,y}^{e})=(l_{1}(\mathbf{\theta},\mathbf{y}^{e}),\cdots,l_{n}(\mathbf{\theta},\mathbf{y}^{e}))^{\prime}$, and $l_{i}(\mathbf{\theta},\mathbf{y}^{e})=\sum_{r=1}^{\infty}\Phi\left(\lambda\mathbf{g}_{i}\mathbf{y}^{e}+\mathbf{z}_{i}^{\prime}\mathbf{\Gamma}-a_{r}\right)$. For any fixed $\bar{R}$, since $\mathbf{y}^{e}$ is not observed, it needs to be computed for every value of $\mathbf{\theta}$. Alternatively, the parameters could be estimated using the NPL algorithm in [42], which takes advantage of an iterative process. The algorithm maximizes a pseudo-likelihood function:
@@
The $\epsilon$ could be replaced by $h_{\epsilon}(\mu_{i},\nu_{i},\bar{\mu}_{i},\bar{\nu}_{i})+\epsilon_{i}^{*}$. Adapting Assumption 2 and Assumption 3 to $\epsilon_{i}^{*}$, the defined Bayesian Nash Equilibrium is still valid. The estimator computes $\hat{\mu}_{i}$ and $\hat{\nu}_{i}$ using a standard Logit in the first step, then substitute the estimated values for $\mu_{i}$ and $nu_{i}$ in the second step. Function $h_{\epsilon}$ is approximated using a sieve method. We refer readers that are interested in the model and technical details to the original paper [34].
The Economics scholar data contains the annual number of publications for each scholar from 2018 to 2021. We split the sample into pre-Covid and Covid periods by time (2018-2019, 2020-2021), and apply the model to each of these 2-year periods to estimate the peer effect. Specifically, We define $y_{i}$ as the total number of publications of scholar $i$ within each 2-year period. Data is segmented in this way because productivity and collaboration are observed to move in different directions before and after 2019. The pandemic and schools’ transition to virtual learning in 2020-2021 also inevitably affect scholars’ productivity and the way they collaborate, so the peer effect $\lambda$ is expected to have some change. We keep the length of each period the same to ensure that the estimated results in the two models are comparable. We include gender, an indicator for African American scholars, expertise, recent productivity indicated by the average number of publications of the scholar in the previous 3 years, total citations up to the first year of each period, and academic experience in the observable characteristics. We discretize scholars’ recent productivity, total citations, and academic experience, and set scholars with an average number of 0 or 1 publication in the previous 3 years, less than 100 citations up to the first year of each period, and less than 10 years of experience, as the reference levels respectively. In the presence of productivity rise during Covid times, we construct an additional Covid index for the period 2019-2021 in order to control for the extent to which the scholars’ publications after the outbreak of the epidemic are related to the Covid context. We use topic modeling to predict the topic of each paper published within these three years based on the title and abstract. The model is capable of recognizing papers that are inspired by the pandemic from other publications, see the table of predicted topics and the words with the highest conditional probability for each topic in the following table. We define a paper is Covid-related if its probability of belonging to the "Covid-19" topic is higher than 50%. The Covid index for each scholar is calculated as:
@@
The mean Covid index for economics scholars is 0.11, meaning that on average, 11% of the papers written in 2019-2021 by an economics faculty are inspired by the pandemic.
@@
To account for the endogeneity of network formation, in the first stage, we estimate scholars’ unobserved characteristics using a dyadic Logit model. In the network formation model, from the school level, we control for homophily of the same department, and the same US News Ranking (Top 10, 11-20, 21-30, 31-40, 41-50). From the individual level, we consider the differences between the two scholars’ academic experience in years, total citations up to the first year of the period, the average number of publications each year during the previous three years, and the number of total publications. We include dummies to indicate whether the dyad involves at least one female author, and at least one African American scholar. We also control for the two scholars’ common research interests, measured by the number of fields that are listed on both scholars’ websites. In the second stage, We include individual fixed effects as additional control variables to evaluate the peer effect.

### v2

# Networks And Productivity – A Study In Economic Scholars During COVID-19

@@ Abstract
The COVID-19 pandemic has disrupted traditional academic collaboration patterns, offering a unique opportunity to analyze the influence of peer effects and collaboration dynamics on research productivity. Using a novel network dataset, this paper investigates the role of peer effects on the productivity of economists, as measured by their publication count, during both pre-pandemic and pandemic periods. The results indicate that peer effects were significant in the pre-pandemic period but not during the pandemic. Additionally, the study sheds light on gender and race differences. These findings enhance our understanding of how research collaboration impacts knowledge production and provide insights that may inform policies aimed at promoting collaboration and boosting research productivity within the academic community.
Keywords Productivity  $\cdot$ Networks  $\cdot$ Network Game  $\cdot$ Peer Effects  $\cdot$ Incomplete Information
@@ 1 Introduction
Academic productivity is a critical metric in decisions related to academic hiring, promotions, and evaluations of research programs and departments (Gingras, 2016; Bornmann et al., 2008). However, measuring academic productivity is a complex question. In general, a higher number of publications is often indicative of a scholar’s active research involvement and significant contributions to the field of study. Recently, there has been an increasing focus on identifying factors that enhance the productivity of scholars across various disciplines. Among these, the role of peer effects and co-authorship has been recognized as significantly influential (Ductor et al., 2014). Scholars with robust social networks and interactions within their field are likely to publish more due to greater access to resources, including research funding, data, and other materials. Co-authorship further promotes knowledge sharing, potentially sparking new research ideas and increasing publication opportunities. Notably, (Petersen, 2015) highlights how a single strong connection can significantly boost a scholar’s productivity and citation rates.
@@
The COVID-19 pandemic has fundamentally altered the way we live, work, and collaborate. As it forced the closure of universities and research institutions, the pandemic inadvertently reshaped the landscape of academic research collaboration. This unprecedented situation offers a unique opportunity to examine the peer effects and co-authorship dynamics among economic scholars during both pre-pandemic and pandemic periods. Working from home has potentially transformed scholars’ approaches to writing, communication, and collaboration. On one hand, the flexibility of schedules and greater autonomy over their work have allowed for enhanced virtual collaboration through video conferencing and instant messaging among scholars located in different regions or time zones. On the other hand, working from home may lead to greater isolation and fewer opportunities for informal interactions and discussions with co-authors and potential co-authors. It may also introduce distractions and disruptions, such as the challenge of balancing work responsibilities with home duties.
@@
This paper utilizes a novel dataset derived from Google Scholar, encompassing scholarly literature and co-authorship networks of economic scholars. We estimate peer effects on the number of publications by these scholars during both pre-pandemic and pandemic periods. The study holds significant implications for several reasons. Firstly, it sheds light on the role of co-authorship in academic productivity, offering insights into how research collaborations contribute to knowledge production and dissemination. Secondly, by contrasting the periods before and during the COVID-19 crisis, this study enhances our understanding of how the pandemic has altered academic collaboration patterns and overall scholarly productivity in economics. Lastly, the findings also illuminate gender and race differences, offering insights that may guide policies aimed at enhancing collaboration and increasing research productivity within the academic community.
@@
This paper is organized as follows: Section 2 reviews the relevant literature on peer effects and academic productivity. Section 3 describes the data and provides summary statistics. Section 4 outlines the model framework and the empirical strategy employed. Section 5 discusses the main results. Finally, Section 6 concludes the paper and suggests directions for future research.
@@ 2 Literature Reivew
From a methodological perspective, our study aligns closely with two main branches of the literature. Firstly, it contributes to the rapidly expanding body of research on peer effects within networks. As illustrated by (Manski, 1993), peer influences manifest in two distinct forms: endogenous peer effects, which pertain to the influence of peers’ outcomes, and contextual peer effects, which relate to the influence of peers’ characteristics. Distinguishing between these effects is challenging due to the simultaneity of behaviors among interacting agents. Pioneering studies on network interactions such as (Bramoullé et al., 2009; De Giorgi et al., 2010; Lin, 2010; Laschever, 2005) lay the groundwork in this field. For instance, (Bramoullé et al., 2009) establishes the benchmark linear-in-means model of peer effects, detailing the identification conditions under the assumption that peers’ influence does not extend beyond direct connections. This assumption—that agents’ characteristics influence individual outcomes solely through their effect on peers’ outcomes—provides valid instruments for addressing correlated effects, thus allowing for the identification of both endogenous and contextual peer effects. A critical insight is that the identification of these effects heavily depends on the network structure. Subsequent research by (Bramoullé, 2013; Arduini et al., 2019a, b; Beugnot et al., 2019) explores heterogeneous peer effects, noting differential impacts among genders. (Masten, 2018) extends this analysis by incorporating heterogeneity, proposing that coefficients for endogenous peer effects vary randomly within a linear-in-means model, and demonstrating that these variations can be precisely identified in the absence of contextual effects for any exogenous characteristic.
@@
The second strand of literature focuses on addressing the issue of correlated effects and leveraging the identification possibilities offered by interaction networks. Researchers have developed at least four broad strategies to tackle this issue: random peers, random shocks, structural endogeneity, and panel data. The first strategy, random peers, involves peers who are randomly allocated through natural or designed experiments. For example, (Sacerdote, 2001) examines roommate pairings at Dartmouth College, (Falk and Ichino, 2006) pairs workers randomly in a laboratory setting, and (De Giorgi et al., 2010) investigates major selections among undergraduates at Bocconi. The key insight here is that with random peers, an agent’s observed and unobserved characteristics are uncorrelated with those of their peers. The second strategy, random shocks, is exemplified by studies like (Dieye et al., 2014), which use exogenous variations for treatment randomization within a linear-in-means framework to identify causal impacts despite potential endogeneity in the network. Other studies, such as (Miguel and Kremer, 2004; Kremer and Miguel, 2007; Crépon et al., 2013), explore spillover effects, highlighting the complexities of estimating causal impacts when outcomes may depend on a comprehensive array of treatments. The third strategy, termed the structural framework, was initially proposed by (Goldsmith-Pinkham and Imbens, 2013) to manage correlated effects by potentially controlling for network endogeneity in peer effect regressions. This approach, reminiscent of Heckman’s correction, aims to account for sample selection and unobserved common factors simultaneously. The fourth strategy pertains to the use of panel data to analyze peer effects, although this aspect of the literature is relatively undeveloped. A few studies, such as (Patnam and Sarkar, 2011; Comola and Prina, 2021; De Giorgi et al., 2020), have employed panel data incorporating individual fixed effects. However, they generally do not tackle contextual peer effects related to time-invariant characteristics, marking a significant limitation and an important area for future research.
@@
The model employed in the present study belongs to the structural endogeneity framework. It is related to models that use static games with incomplete information, in which agents act non-cooperatively (see (Harsanyi, 1967; Osborne and Rubinstein, 1994)). The assumption of incomplete information in peer effect models for discrete outcomes is widely studied (e.g., (Brock and Durlauf, 2001; Bajari et al., 2010; Yang and Lee, 2017; De Paula, 2017)). In the literature, an agent $i$’s decision is influenced by their own observable characteristics, unobservable individual type, and other agents’ choices. Recent studies by (Boucher and Houndetoungan, 2020; Houndetoungan, 2022) propose methods to estimate the network’s probability distribution using cross-sectional data when the network is imperfectly observed. They construct a network game where each agent chooses an integer outcome to maximize their preference, which includes observed characteristics of the agent and peers, the difference between the agent’s choice and that of peers, a cost function, and a private signal. They prove that under certain assumptions, there is a unique Bayesian Nash Equilibrium for this game, and they propose an estimator based on pseudo-likelihood maximization.
@@
The literature of empirical papers exploring the relationship between co-authorship and academic productivity has expanded recently. Nonetheless, consensus remains elusive regarding whether this relationship is positive, negative, or insignificant. For example, (Cainelli et al., 2015) demonstrates that economists who are more collaborative are also more productive, and factors such as tenure, age, and geographical variables do not have a significant impact on productivity. Similarly, (Ductor, 2015) finds a positive correlation between intellectual collaboration and individual performance after accounting for endogenous network formation, unobservable heterogeneity, and factors that vary over time. As indirect evidence, (Bosquet and Combes, 2013) identifies that, at the individual level, the average publication quality rises with the average number of authors per paper, individual field diversity, the total number of published papers, and the presence of foreign co-authors, while female and older academics tend to publish less frequently.
Conversely, some scholars argue that the relationship between co-authorship and productivity exhibits a negative correlation or lacks statistical significance. (Hollis, 2001) discovers that for a specific scholar, increased co-authorship leads to higher quality, longer, and more frequent publications. However, after adjusting for the number of authors, the relationship between co-authorship and a scholar’s attributable output becomes negative. Other research contends that after controlling for article length, journal and author quality, subject area, and scholar fixed effects, the productivity of prior collaborators is not a significant determinant of a researcher’s own productivity (Cheng, 2022) or higher quality research (Medoff, 2003). Furthermore, (Oettl, 2012) finds that co-authors of highly helpful scientists who die experience a decrease in output quality but not output quantity.
@@ 3 Economic Scholars Data
The Economic Scholars dataset consists of 1,671 core faculty members from the top 50 Economics Schools in the United States, based on the 2022 US News rankings, who have registered a homepage on Google Scholar. The dataset excludes visiting professors, teaching professors, and lecturers. These homepages provide rich information on the individuals’ influences and research journeys in academia. Besides the names and affiliations, scholars may list their research interests and sub-fields at the top of the page. On the right-hand side, the scholar’s cumulative citations, H-index, and I10-index are displayed, along with regular coauthors and a histogram showing the number of citations and publications each year for the last 10 years. On the left-hand side, a comprehensive list of the scholar’s academic output can be viewed, including the paper’s title, journal, authors, year, and the number of citations. See a snapshot of the Google Scholar homepage in the figure below.
@@
A key feature of the data collected from Google Scholar is its comprehensiveness. The website aggregates information from a wide range of sources, including journal articles, conference papers, theses and dissertations, technical reports, books and book chapters, patents, working papers, and repositories. Thus, compared to using data sources from academic journals, the lagged effect of co-authorship on scholars’ academic output due to journals’ reviewing processes can be largely alleviated.
Based on scholars’ publications, we construct an economics co-authorship network for each year over the last 10 years. Each node represents a faculty member in the economics department, and there is an edge between any two nodes if they have co-authored at least one paper in a given year. The edge density of a temporal network indicates the prevalence of scholar collaborations in that year, or the percentage of observed co-authorships over all possible collaborations between any two scholars. Economic scholars’ productivity, as revealed by the annual average number of publications, gradually increased from 2013, surged to a peak in 2020, and then slowly declined in 2021 and 2022. Network edge density reached its highest level in 2018, declined to its lowest point in 2020, and then rebounded in 2021 and 2022. Interestingly, scholars’ average productivity moved in the same direction as the prevalence of collaborations before 2018 but in the opposite direction from 2019 to 2021.
@@
The experience of economic faculty in academia averages 25 years as of 2023, measured by the difference between 2022 and the year of the scholar’s first published article on Google Scholar. African American scholars, identified either directly by their listed nationality in CVs or indirectly by the combination of the predicted country of origin using names and faculty pictures, account for 0.9% of the economics faculty. In our sample, 18% of the faculty are female.
We also label the expertise of economic scholars with their listed sub-fields on their Google Scholar homepage. For those who have left this section blank, we fill in the missing values using the fields shown on the department faculty website and the scholar’s curriculum vitae. The sub-fields are not mutually exclusive, and it is reasonable for scholars to have multiple specialties. For example, a macroeconomist may specify econometrics, monetary policy, international economics, or finance as their expertise. In our sample, economic/econometric theory, macroeconomics, and labor economics are the three top fields that scholars work on, followed by econometrics, industrial organization, and development economics.
Table 1: Summary Statistics of Economic Scholars
@@
Table 2: Sub-fields of Economic Scholars
@@ 4 Network Game with Peer Effects and Incomplete Information
Following the notation in (Houndetoungan, 2022), suppose there is a population of $n$ agents that interact through a network matrix $\mathbf{G}$ with zero diagonal and non-negative elements $g_{ij}$ that represents the proximity of $i$ and $j$. Each agent $i$ chooses an integer outcome, $y_{i}$, to maximize his or her individual utility:
@@
Define $\Delta$ as the first difference operator. It is proved in (Houndetoungan, 2022) that under the following three assumptions, there is a unique integer $r_{0}$ that maximizes the preference $U_{i}^{e}\left(\cdot\right)$, and $U_{i}^{e}\left(r\right)\geq\max\left\{U_{i}^{e}\left(r-1\right),U_{i}^{e}\left(r+1\right)\right\}$ if and only if $r=r_{0}$:
@@
where $\Phi(\cdot)$ is the CDF of standard normal distribution, $\mathbf{L}$ is a mapping, $\mathbf{y}^{e}=\mathbf{L}(\mathbf{\theta,y}^{e})=(l_{1}(\mathbf{\theta},\mathbf{y}^{e}),\cdots,l_{n}(\mathbf{\theta},\mathbf{y}^{e}))^{\prime}$, and $l_{i}(\mathbf{\theta},\mathbf{y}^{e})=\sum_{r=1}^{\infty}\Phi\left(\lambda\mathbf{g}_{i}\mathbf{y}^{e}+\mathbf{z}_{i}^{\prime}\mathbf{\Gamma}-a_{r}\right)$. For any fixed $\bar{R}$, since $\mathbf{y}^{e}$ is not observed, it needs to be computed for every value of $\mathbf{\theta}$. Alternatively, the parameters could be estimated using the nested pseudo likelihood (NPL) algorithm in (Aguirregabiria and Mira, 2007), which takes advantage of an iterative process. The algorithm maximizes a pseudo-likelihood function:
@@
The $\epsilon$ could be replaced by $h_{\epsilon}(\mu_{i},\nu_{i},\bar{\mu}_{i},\bar{\nu}_{i})+\epsilon_{i}^{*}$. Adapting Assumption 2 and Assumption 3 to $\epsilon_{i}^{*}$, the defined Bayesian Nash Equilibrium is still valid. The estimator computes $\hat{\mu}_{i}$ and $\hat{\nu}_{i}$ using a standard Logit in the first step, then substitute the estimated values for $\mu_{i}$ and $nu_{i}$ in the second step. Function $h_{\epsilon}$ is approximated using a sieve method. We refer readers that are interested in the model and technical details to the original paper (Houndetoungan, 2022).
The Economics scholar data contains the annual number of publications for each scholar from 2018 to 2021. We split the sample into pre-Covid and Covid periods (2018-2019 and 2020-2021) and apply the model to each of these 2-year periods to estimate the peer effect. Specifically, We define $y_{i}$ as the total number of publications of scholar $i$ within each 2-year period. Data is segmented in this way because productivity and collaboration are observed to move in different directions before and after 2019. The pandemic and schools’ transition to virtual learning in 2020-2021 also inevitably affected scholars’ productivity and collaboration, so the peer effect $\lambda$ is expected to have some change. We keep the length of each period the same to ensure that the estimated results in the two models are comparable.
We include gender, an indicator for African American scholars, expertise, recent productivity indicated by the average number of publications of the scholar in the previous 3 years, total citations up to the first year of each period, and academic experience in the observable characteristics. We discretize scholars’ recent productivity, total citations, and academic experience, setting scholars with an average of 0 or 1 publication in the previous 3 years, fewer than 100 citations up to the first year of each period, and less than 10 years of experience as the reference levels respectively.
In the presence of a productivity rise during Covid times, we construct an additional Covid index for the period 2019-2021 to control for the extent to which scholars’ publications after the outbreak of the epidemic are related to the Covid context. We use topic modeling to predict the topic of each paper published within these three years based on the title and abstract. The model is capable of distinguishing Covid-inspired papers from other publications. See the table of predicted topics and the words with the highest conditional probability for each topic in the following table. We define a paper as Covid-related if its probability of belonging to the "Covid-19" topic is higher than 50%. The Covid index for each scholar is calculated as:
@@
The mean Covid index for economic scholars is 0.11, meaning that on average, 11% of the papers written by an economics faculty member in 2019-2021 are inspired by the pandemic.
@@
To account for the endogeneity of network formation, in the first stage, we estimate scholars’ unobserved characteristics using a dyadic Logit model. In the network formation model, at the school level, we control for homophily by considering whether scholars are from the same department and the same US News Ranking category (Top 10, 11-20, 21-30, 31-40, 41-50). At the individual level, we consider differences between the two scholars in terms of academic experience in years, total citations up to the first year of the period, average number of publications each year during the previous three years, and total number of publications. We include dummies to indicate whether the dyad involves at least one female author and whether it involves at least one African American scholar. We also control for the two scholars’ common research interests, measured by the number of fields listed on both scholars’ websites. In the second stage, we include individual fixed effects as additional control variables to evaluate the peer effect.